OpenStudy (anonymous):

if y=16 when x=4 find y when x=6

OpenStudy (anonymous):

help

OpenStudy (queelius):

This is a very odd question. I suspect you did not post the complete question. The only thing we know about the function f is that it maps x to 16 -- that is, y = f(4) = 16. However, we do not have any other information about f. There are an uncountably infinite number of other functions that map 4 to 16, e.g., f(x) = 12 + x or f(x) = x^2.